OpenStudy (anonymous):

Point Slope Form of the equation of the line Slope=7 Passing through -6,5

OpenStudy (anonymous):

ok so we know that y=mx+ b where m is slope. we just plug in your given values so 5=(7)(-6)+ b. from there we see that b is 5+42 which is 47. so from there we just write it in point slope form which is y=7x+47

jimthompson5910 (jim_thompson5910):

y - y1 = m(x - x1) y - 5 = 7(x - (-6)) y - 5 = 7(x + 6)